Complete the statement using the
a)Distributive property
b)Commutative property of addition
c)Commutative property of multiplication
6 ∙ (5 + 4) =




Answer by scianci(186)

a. 6(5 + 4) = 6(5) + 6(4)
b. 6(5 + 4) = 6(4 + 5)
c. 6(5 + 4) = (5 + 4)(6)